My problem is to solve 2- (a-4/a+4)=(a^2-8/a+4)I think I need to multiply each side of the equation by a+4 to eliminate fractions. So I end up with (a+4)*2-(a-4)=a^2-8. That's as far as I've gotten. I would really appreciate any help to complete the problem. Thank you.
Answer by edjones(8007)

2- (a-4/a+4)=(a^2-8/a+4)
2(a+4)-(a-4)=a^2-8 multiply each side by a+4
2a+8-a+4=a^2-8
a+12=a^2-8
a^2-a-20=0 Subtract the left side from the right.
(a-5)(a+4)=0
a=5 (a cannot equal -4 because there would be division by zero which is not allowed).
.
Ed